OpenStudy (anonymous):

Given a = 4, b = -2 and c = 8, evaluate a + b + c.

OpenStudy (anonymous):

a + b + c = 4 -2 +8

OpenStudy (anonymous):

can you evaluate the final answer???

OpenStudy (anonymous):

yes (: Can You Help With More

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

Given a = 4, b = -2 and c = 8, evaluate ab - bc

OpenStudy (anonymous):

ab - bc = (4)(-2) = (-2)(8)

OpenStudy (anonymous):

Given a = 4, b = -2 and c = 8, evaluate 3abc

OpenStudy (anonymous):

3abc = 3(4)(-2)(8)

OpenStudy (anonymous):

Given a = 4 and b = -2, evaluate -|b - a|.

OpenStudy (anonymous):

-|b - a| = -|-2 - 4| = -|-6| = -6

OpenStudy (anonymous):

3(2x - 4) = 4(x + 1)

OpenStudy (anonymous):

3(2x - 4) = 4(x + 1) 6x - 12 = 4x +4 6x - 4x = 12+4 2x = 16 x = 8

OpenStudy (anonymous):

|3 - x| = 8 , I need As Much help As I Can I Graduate Next Week And Its Confusing Me + Im Behind